OpenStudy (ttop0816):

Which of the following ordered pairs lies on the graph of y = tanx? (-5pi/2, -1) (-9pi/4, 1) (5π, 0)

OpenStudy (zzr0ck3r):

what is \(\tan(\frac{-5\pi}{2})\)?

OpenStudy (zzr0ck3r):

@ttop0816 what is \[\frac{\sin(\frac{-5\pi}{2})}{\cos(\frac{-5\pi}{2})}\]

OpenStudy (zzr0ck3r):

do you know \[\sin(\frac{\pi}{2})\]?

OpenStudy (ttop0816):

is it 1??

OpenStudy (zzr0ck3r):

do you know \(\sin(-\frac{\pi}{2})\)?

OpenStudy (zzr0ck3r):

correct

OpenStudy (ttop0816):

and sin (-pi/2) would be -1 then?

OpenStudy (zzr0ck3r):

so what is sin(-5pi/2)?

OpenStudy (ttop0816):

-1!

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

what is cos(-5pi/2)

OpenStudy (ttop0816):

0

OpenStudy (ttop0816):

oh then for each (-9pi/4, 1) & (5π, 0) i have to do for example sin (-9pi/4) & cos (-9pi/4) ... more??

OpenStudy (zzr0ck3r):

so \(-1\ne\tan(-\frac{5\pi}{2})\) so (-5pi/2,-1) is not a solution

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

one of those will work out

OpenStudy (zzr0ck3r):

well notice we are looking at \[\tan(x) = \frac{sin(x)}{cos(x)}\]

OpenStudy (ttop0816):

ohhhhh!!!! then wait

OpenStudy (zzr0ck3r):

so for -pi/4 we do \[\frac{sin(-pi/4)}{cos(-pi/4)}\]and see if that equals 1 if it does, then its a solution

OpenStudy (ttop0816):

i got (-9pi/4, 1)

OpenStudy (zzr0ck3r):

sin(-9pi/2) = -sqrt(2)/2 cos(-pi/2) = sqrt(2)/2 divide them and you get -1

OpenStudy (ttop0816):

oh and where did you get -pi/2 from?? also would i use the x cord to calculate sin (-9pi/4) and use the y cord to solve cos (1)?

OpenStudy (ttop0816):

am i correct with my answer overall?